OpenStudy (anonymous):

Determine if fifth degree polynomials are a vector space? Do they fit the requirements

OpenStudy (anonymous):

James I proved the rest of the identities on my own LOL :D

OpenStudy (amistre64):

what are the requirements?

OpenStudy (jamesj):

No. For example, here are two fifth degree polynomials : v1(x) = x^5 + x^4 and v2(x) = -x^5, But v1(x) + v2(x) = x^4, which is not a fifth degree polynomial.

OpenStudy (anonymous):

ohhh i see

OpenStudy (jamesj):

That being said, the set \( P_5 \) being the defined as the set of all polynomials \( \it up \ to \) fifth degree is a vector space.

OpenStudy (amistre64):

ive been wondering if James has a pool of knowledge floating around inside his skull; or if its really a rolodex next to the computer with all this stuff in it :)

OpenStudy (anonymous):

heheehehe

OpenStudy (jamesj):

I bought it off a Tibetan voodoo monk: "The Book of Answers". Cost me fifty gold dragoons and the blood of a che-wang white bat. But it was worth it.

OpenStudy (amistre64):

whats on page 42?

OpenStudy (jamesj):

the multiplication table for for 7.

OpenStudy (amistre64):

I must have a later edition then :/ mine says: *

OpenStudy (anonymous):

LOL u guys r funny Thanks for ur help :D

OpenStudy (jamesj):

When I say the multiplication table for 7, I mean the whole thing. It's in infinitely small type. Can't read a damn thing.